# Lasso Multinomial Performance Indicators for in-play Basketball Data

### v1

@@ 5 Highlights and Discussion
Finally, considering insights from the multinomial model, we introduce a novel RAPM metric: the Expected Points (EPTS). This metric is derived from a correctly specified model for the number of points scored. To further refine the analysis, we propose a weighted version of EPTS, denoted as wEPTS. This extension incorporates player participation within each team’s possessions, effectively handling the issue of inflated contributions observed with low-time players. In summary, EPTS is based on a statistically sound foundation for evaluating player performance based on expected points scored. Moreover, our evaluation using established validation criteria demonstrates that wEPTS outperforms all other approaches considered in this study.
@@
To conclude, in this study, we have considered a variety of issues concerning the estimation of RAPM player ratings from possession-based NBA data. The main contributions of this paper are: (a) we introduce the use of lasso instead of ridge and demonstrate its superiority in terms of evaluations; (b) we offer a clear interpretation of commonly used normal ridge RAPMs through logistic regression coefficients; and finally, (c) we introduce novel RAPM metrics: Expected Points Scored (EPTS) and its weighted version (wEPTS) based on a multinomial logistic regression model, which is the statistically appropriate approach for modelling the points scored per possession. The proposed wEPTS has four clear advantages: First, it outperforms all other approaches evaluated in this study, including the standard ridge normal RAPMs. Secondly, it effectively addresses the challenge of low-time players, almost eliminating their inflated contributions. Third, it is based on a foundationally appropriate statistical model and, four, it can separately evaluate and consider the contribution of each player based on the different type of scoring points that each player contributes.

### v2

@@ 6.1 Conclusion
Considering insights from the multinomial model, we introduce a novel RAPM metric: the Expected Points (EPTS). This metric is derived from a correctly specified model for the number of points scored. To further refine the analysis, we propose a weighted version of EPTS, denoted as wEPTS. This extension incorporates player participation within each team’s possessions, effectively handling the issue of inflated contributions observed with low-time players. In summary, EPTS is based on a statistically sound foundation for evaluating player performance based on expected points scored. Moreover, our evaluation using established validation criteria demonstrates that wEPTS outperforms all other approaches considered in this study.
Considering our reviewers’ suggestions, and given the discrete or ordinal nature of points per possession (our response), two additional models were examined; a Poisson and an ordered multinomial regression model. The results confirmed our initial intuition that these models were not appropriate for our response data (i.e. the points per possession). Firstly, the Poisson model incorrectly assumes that the probability of scoring one point is higher than two and three, which is not generally true in basketball. Secondly, the proportional odds assumption for the ordered multinomial model is violated, since the effect of a player on the odds of one scoring category compared to all higher categories is not constant across all thresholds of the ordinal outcome.
@@
To conclude, in this study, we have considered a variety of issues concerning the estimation of RAPM player ratings from possession-based NBA data. The main contributions of this paper are: (a) we suggest the use of lasso instead of ridge and demonstrate its superiority in terms of evaluations; (b) we offer a clear interpretation of commonly used normal ridge RAPMs through logistic regression coefficients; and finally, (c) we introduce novel RAPM metrics: Expected Points Scored (EPTS) and its weighted version (wEPTS) based on a multinomial logistic regression model, which is the statistically appropriate approach for modelling the points scored per possession. The proposed wEPTS has four clear advantages: First, it outperforms all other approaches evaluated in this study, including the standard ridge normal RAPMs. Secondly, it effectively addresses the challenge of low-time players, almost eliminating their inflated contributions. Third, it is based on a foundationally appropriate statistical model and, four, it can separately evaluate and consider the contribution of each player based on the different type of scoring points that each player contributes.
6.2 Discussion and future research
Due to the nature of basketball, where players interact in offence and defence in order to perform better as a team, analyzing the impact of lineups, rather than solely individual players, is quite important. The Regularized Adjusted Plus-Minus (RAPM) evaluates players performance, while their teammates and opponents are considered in its development. The weighted Expected Points per Possession (wEPTS) we propose in this work has a similar value, but at the same time, it addresses issues found in Gaussian RAPMs, such as accounting for the discrete nature of the response variable, the varying contributions of players to different types of points, and the effect of LTPs.
The proposed rating of weighted EPTS (wEPTS) appears with approximately 78% agreement (while 63% corresponds to EPTS) between the lineups determined for each team, considering the highest offensive and defensive contributed players per position, and the two most frequently used (most-played) lineups per team in the season. Although this result seems surprising, it does not imply that simply combining individual player contributions would lead to a useful evaluation or construction of more effective lineups, since our approach is designed to capture the individual players’ effects.
In the literature, the effect of lineups is examined using methodologies similar to those applied to individual player performance ratings, replacing individual players with lineups in the analysis. This simplified solution could be applied to estimate the lineups effect through a multinomial logistic model.
Regarding this, [9] introduced an alternative way to estimate lineups effect, including both lineup and players effects in the model specification; suggesting that the first can be seen as a type of higher-order interaction among players. In fact, each player demonstrates a varying level of impact depending on the lineup in which he is included. Similar ideas could be introduced in our proposed methodology, by incorporating the effect of lineups or similar adjustments in order to capture or at least consider the interaction.
In this context, “interaction” is the key, since investigating and detecting possible patterns in the court is more than desirable. Within a lineup, each player has a specific role that complements the roles of their teammates. Therefore, considering the abilities and tendencies of each player, coaches and staff try to optimize their roster under dynamic situations. This is the main challenge of such analyses; dynamic situations arise from a variety of factors. Thus, an analysis grounded in a team’s in-game habits and style of play may prove to be both promising and highly valuable, but also challenging.
Considering the disaggregated data that were used, some relevant covariates may be useful. For instance, [9] included the game period and the type of shots (close-range, mid-range, three-pointers) in their proposed model, while they used some boxscore statistics to define the response variable based on the outcome of a possession. Following these ideas, some candidate covariates would be the time till the end of the game (the garbage time could be considered as an extra indicator), the score difference at each possession-event (or a variation according to how a field goal made decreases or increases the score difference of a team), the time of an attempted shot (made or not) or a turnover during offence (fast break, secondary transition, and, set), which can be connected to a team’s style of play.
Moreover, the competition level and structure of each basketball league vary. For instance, an average NBA team has the opportunity to compete for postseason qualification even in the most recent regular season games, but an average Euroleague team needs to try harder during the regular season in order to proceed to the playoffs. This aspect could be reflected through the importance of a game within the season and mathematically converted into an appropriate, and dynamic weight.
Thus, in order to improve the accuracy of the players’ performance ratings, the proposed approach can include the extra variables described in the previous two paragraphs.
Moreover, the suggested methodology might incorporate different shrinkage techniques. Avoiding underestimation or overestimation of players contributions, especially in the case of low-time players, is of main interest. In recent years, there has been a great enlargement in shrinkage literature, especially through Bayesian techniques. Global-local shrinkage methods, like spike-and-slab (introduced by 5) and horseshoe (introduced by 2) prior assuming on players effect seem suitable and promising for our case study. These two approaches are similar; in the first, the local shrinkage parameter is typically fixed, whereas in the second, it is distributed using a half-Cauchy distribution. As an additional step in our suggested process, we can take player participation into account for this parameter. Alternatively, we may choose not to apply any of these shrinkage priors and instead set an “informative” prior based on the playing time of each player to account for their individual effect. Even while this sounds interesting, we need to take certain computational issues into consideration.